I have a right-angle triangle with a hypotenuse of 4.6cm and an area of 2cm. I need to calculate the lengths of the other two sides. I can see the product of the unknown sides is 4, but is there a formula to calculate the ratio of the sides and therefore the lengths?
Answer by richwmiller(17219)

1/2 bh=2
b*h=4,
b^2+h^2=4.6^2= 21.16
b=4.51384
h=.886164